OpenStudy (anonymous):

PLEASE HELP!!! A recursive function is shown below: f(1) = 2 and f(n) = f(n - 1) - 5; n > 1 Which of the following lists the terms in the sequence defined by this recursive function? 2, 3, 8, 13, 18, ... 2, 7, 12, 17, 22, ... 2, -3, -8, -13, -18, ... 2, -7, -12, -17, -22, ...

OpenStudy (anonymous):

@hartnn

OpenStudy (zzr0ck3r):

just plug things in

OpenStudy (anonymous):

Is it C?

OpenStudy (zzr0ck3r):

f(1) = 2 and f(2) = f(2-1) - 5= f(1) - 5 = 2-5 = -3 so C correct

OpenStudy (anonymous):

thanks :)